OpenStudy (anonymous):

HELP ME (my BOO's<3) 30% off of an item that coasts $89.00

jimthompson5910 (jim_thompson5910):

30% of 89 = 0.3*89 = 26.7 So 30% off of 89 is 89 - 26.7 = 62.3

jimthompson5910 (jim_thompson5910):

Alternatively, 30% off of 89 = 70% of 89 = 0.7*89 = 62.3